# PaECTER: Patent-level Representation Learning using Citation-informed Transformers

### v1

@@ Abstract
PaECTER is a publicly available, open-source document-level encoder specific for patents. We fine-tune BERT for Patents with examiner-added citation information to generate numerical representations for patent documents. PaECTER performs better in similarity tasks than current state-of-the-art models used in the patent domain. More specifically, our model outperforms the next-best patent specific pre-trained language model (BERT for Patents) on our patent citation prediction test dataset on two different rank evaluation metrics. PaECTER predicts at least one most similar patent at a rank of 1.32 on average when compared against 25 irrelevant patents. Numerical representations generated by PaECTER from patent text can be used for downstream tasks such as classification, tracing knowledge flows, or semantic similarity search. Semantic similarity search is especially relevant in the context of prior art search for both inventors and patent examiners. PaECTER is available on Hugging Face. [^1] https://huggingface.co/mpi-inno-comp/paecter
@@ 1 Introduction
Domain-specific language models have become increasingly vital in social sciences and technology, transforming text into valuable data. Social scientists, in particular, have shown a keen interest in conducting similarity comparisons between patents. Existing state-of-the-art tools are tailored explicitly to scientific English, which may differ from patent-specific language. The PaECTER (Patent Embeddings using Citation-informed TransformERs) encodes patents into numerical representations to detect semantic similarity.
@@
The introduction of the Bi-Directional Encoder Representations from Transformers (BERT) by Devlin et al., (2019) was a major advancement in the toolbox of scientists working with text. These kinds of models are trained on huge textual corpora and generate fixed-size dense vectors (embeddings) from text. BERT and its derivatives work with a fixed vocabulary of so-called tokens (words or sub-words). Tokens not found in this vocabulary are either discarded or fragmented into smaller, recognizable tokens. A set of tokens may not adequately encompass the complete spectrum of linguistic features found in particular language usage, leading to the creation of language-specific BERT derivatives.
@@
Therefore, researchers have developed various BERT models that have been trained on and for specific corpora. Some of them are suitable for the task of transforming patent texts into embeddings, namely BERT for Patents (Srebrovic and Yonamine,, 2020) or SciBERT (Beltagy et al.,, 2019), a BERT-based pre-trained language model specific to scientific language. Having the right vocabulary is crucial for domain-specific tasks. Our analysis reveals that only approximately 50% of the vocabulary used in SciBERT directly overlaps with that of BERT for Patents.
However, since researchers are interested in the similarity between texts, such models have a drawback. They are not ideally suited for such a similarity task, as their training goal was to predict masked tokens and next sentences, not similar documents.
Therefore, a suitable relatedness/similarity signal and a similarity training objective must be identified. To accomplish such a task, contrastive learning is used. This machine learning method finds patterns by teaching the model to recognize similarities and differences between data points during the training. The key element is the triplet loss function and specially crafted training data that contain signals for similarity and dissimilarity.
@@
SPECTER (Cohan et al.,, 2020)and SPECTER2 (Singh et al.,, 2022) use this approach. They make use of the existing capabilities to model scientific language by taking SciBERT as a foundation and generate a training dataset with so-called triplets. These triplets consist of a query document, an actual (positive) citation, and a non-cited (negative) document. The SPECTER model initially used 684,000 training triplets (146,000 query documents), while its extended version, SPECTER2, was trained on over 6 million triplets. By encoding titles and abstracts of scientific publications, SPECTER and SPECTER2 have shown improved performance on multiple downstream tasks in the scientific domain, such as topic classification, citation prediction, and recommendation.
@@
SEARCHFORMER (Vowinckel and Hähnke,, 2023) uses the same approach in the patent domain. However, this model is unsuitable for comparing patent titles and abstracts because of its focus is on the patent claims. Furthermore, it is not publicly available.
This leaves a gap in the patent space for a model with the appropriate vocabulary/language and citation-induced similarity that is publicly available at the same time.
We fill this gap by following the contrastive learning technique used in the previously described models. The foundation of our model is the BERT for Patents, whose learning training objective is to find the semantic similarity between the cited patent and the focal patent with respect to the non-cited patent.
Since we rely on citations as similarity signal, we focus on EPO patents, for two reasons. At the EPO, references references in search reports are under the full control of examiners, mitigating any strategic bias typically associated with citations made by inventors or patent attorneys. Moreover, these references are classified with a unique emphasis on their contribution to the patent filed (Webb et al.,, 2005), a feature not present in US patents. In total, we use 300.000 patent families from the EPO, and generate five triples for each.
We have fine-tuned the SPECTER 2.0 model with our training patent dataset for benchmarking. We have named this model PAT SPECTER. Figure 1, shows how the different models relate to each other.
Figure 1: Overview of the different models
Both models, the PaECTER and the PAT SPECTER can be used for downstream tasks in the patent domain, such as topic classification, citation prediction, and recommendation.
@@ 2 Training Data
Our training data involves 300,000 English-language patent families, including applications filed with the European Patent Office (EPO) from 1985 to 2022. A patent family refers to those patent applications and grants which originate from the same set of priority filings, but seek protection in different jurisdictions.
We specifically focus on EPO patents for two critical reasons. First, references in EPO search reports are under the full control of examiners, mitigating any strategic bias typically associated with citations made by inventors or patent attorneys. Second, these patents are classified with a unique emphasis on their contribution to the patent filed (Webb et al.,, 2005), a feature not present in US patents.
@@ 2.1 Objective
The primary goal of our model is to grasp an understanding of similarity and dissimilarity based on textual content. By feeding the model examples of similar and dissimilar documents in relation to a focal document, the model autonomously learns optimal weights that minimize the loss function. Triplet margin loss learning is an effective approach in this context, where the adjustments of the weights are guided by maximizing the differences between similar and dissimilar examples.
Our training dataset comprises triplets of patents with a focal patent, a similar patent (positive), and a dissimilar patent (negative). Similar patents are those that the focal patent cites directly with a specific citation category. Dissimilar patents are those it does not cite, subject to further refinements.
The triplet margin loss function is obtained as:
@@
where .$\lVert.\rVert_{2}$ represents the $L_{2}$ norm distance, and $V_{F},V_{P},V_{N}$ are numerical representations for the focal patent ($P_{F}$), positive patent ($P_{P}$), and negative patent ($P_{N}$), respectively. The margin variable ($m$) ensures that the positive patent with respect to the negative patent is always closer to the focal patent by at least a certain value.
@@ 2.2 Focal Patents
We selected 300,000 patent applications as focal patents from the PATSTAT 2023 Spring version following specific criteria:
@@ 2.3 Positive Citations
Positive citations are selected exhaustively with replacement from the pool of eligible patent documents for each focal patent, based on specific citation categories (Table 1). Categories X, Y, I, and A are pertinent, indicating a strong similarity or relevance to the state of the art.
@@
X
Particularly relevant if taken alone - prejudicing novelty.
I
Particularly relevant if taken alone - prejudicing inventive step.
Y
Particularly relevant if combined with another document of the same category.
A
Documents defining the state of the art and not prejudicing novelty or inventive step.
O
Non-written disclosure.
P
Intermediate document.
T
Theory or principle underlying the invention.
E
Earlier patent application, but published after the filing date of the application searched (potentially conflicting patent documents).
D
Document cited in the application.
L
Document cited for other reasons.
Member of the same patent family, corresponding document
Source From Webb et al., (2005). Categories in bold font are the ones we are using.
Of these categories, X, Y, I and A are relevant for the training since the focal patents citing the patents of categories X, Y, and I have strong similarity with the cited ones. The cited patent of kind X can even prejudice the novelty of the focal patents. The cited patent of kind A also reflects the state of the art for the focal patents. Thereby, we consider the cited patents of kinds X, Y, I, and A to be similar to the focal patents.
@@ 2.4 Negative Citations
Negative patents, representing dissimilarity with the focal patents, are selected with a nuanced approach. First, they must come from the universe of all patents ever cited by any EPO patent, which was filed during the 1985-2022 period. Second, at least the abstract is available in English through the cited patent itself or its best substitute. Third, a CPC category must exist.
@@
  • Easy Negatives: Uncited patents in the same CPC class and published within five years prior to the focal patent, and not indirectly cited.
@@
The selection process is illustrated in Figure 2. The number of all potential easy negatives is 6,484,216; there was no focal patent that did not meet the requirements of at least three easy negatives.
@@ 2.5 Test Dataset
For the test dataset, we construct 1000 samples consisting of 1 focal patent with 5 positive citations and 25 negative citations. The negative citations are a random sample of all easy negatives of a patent. Since the dataset is smaller, only focal patents were selected where there were enough positive and negative examples. We do not sample with replacement.
@@ 3 Training
We build our model utilizing sentence transformers architecture (Reimers and Gurevych,, 2019) and HuggingFace framework (Wolf et al.,, 2020). At the core of which, we use BERT for Patents (Srebrovic and Yonamine,, 2020) that is connected to an input module at the preceding layer and also connected to a document level embedding module at the following layer.
The input module concatenates a patent’s title and abstract and passes the concatenated text to the BERT for Patents model. Because of the architecture of BERT models, only the first 512 tokens (about 393 words) are used. All the 1024 dimensional tokens of the output layer of the BERT for Patents model subsequently go to the document level embedding module, which produces the mean embedding of all these tokens as the final numerical representation for the given input patent text. [^3] We experimented with the mean embedding of the tokens of the output layer of BERT for Patents model; it outperformed the embedding of [CLS] token in the similarity learning task. Hence, we do not use [CLS] token as the final representation of any given text. During training, we feed each triplet to the model. It then optimizes the triplet loss function (eq. (1)) by comparing the numerical representations of all three documents in that triplet.
We do minimal hyperparameter-tuning for training, following the leads of SPECTER and SEARCHFORMER. Post-tuning, we use learning rate of $1e$-$5$, 10% of the training data for warmup steps, and triplet loss margin ($m$) of 1 for training in addition to the optimizer, Decoupled Weight Decay Regularization (Loshchilov and Hutter,, 2019)
We train our model for 4 epochs, but the training doesn’t improve after the second epoch. Each training epoch spans approximately 20 hours, with validation occurring every 4000 steps.
We use a batch size of 4 (maximum GPU capacity) and gradient accumulation of 8 on four NVIDIA A100-SXM4-40GB GPUs for faster training and achieving an effective batch size of 128. To train the model with multiple GPUs, we have adapted the code base of the sentence transformers package slightly [^4] See the forked GitHub repository:
https://github.com/max-planck-innovation-competition/sentence-transformers
, and also utilized Hugging Face’s Accelerate library.
@@ 4.1 Performance Evaluation
We conduct a rank-aware evaluation of our trained model, PaECTER and compare it against several other models and algorithms on our test dataset. The baseline of our evaluation is the BM25 algorithm. We then compare the performance of SPECTER, SPECTER version 2 (BASE), and BERT for Patents. [^5] We do not include SEARCHFORMER because the model is not available publicly.
Furthermore, we also fine-tuned a variant of SPECTER 2 (BASE) on our patent training data and refer to it as PAT SPECTER. The primary objective of including PAT SPECTER in the comparison is to benchmark PaECTER against a model originating from scientific domain but fine-tuned on our specific domain of interest.
We use three evaluation metrics. The first Rank First Relevant (RFR), which is simply the position of the first actual reference patent within the ranked list of related patents for each sample. Secondly, we use the Mean Average Precision (MAP) assesses the average precision of correctly identifying the 5 actual reference patents within the ranked list of 30 candidate patents for each sample, and then averages these precision scores across all samples. Finally, we use Mean Reciprocal Rank @ 10 (MRR@10), which averages the reciprocal ranks of the first correctly predicted patent among the set of 10 out of 30 best ranked patents, averaged over all test samples.
Moreover, since pooling strategies can potentially affect model’s performance, we consider two widely used pooling strategy during evaluation, such as CLS and mean. Table 2 presents the performances of all models under consideration for both the pooling strategies.
In Table 2, all three metrics point to the superiority of the PaECTER model regardless of which pooling strategy between CLS and mean is used. In particular, our trained model with mean pooling achieves higher accuracy over BERT for Patents by $7.85$ points (significant, p<.001) in MAP and $7.76$ points (significant, p<.001) in MRR@10. Furthermore, Figure 3 illustrates that in terms of RFR score, our model consistently predicts first relevant one within the topmost positions of the predicted outcome compared to other models.
@@
BERT-large-uncased 2.29 1.98 38.17 51.69 58.95 71.06
SciBERT-scivocab-uncased 3.07 1.91 37.76 51.31 58.80 73.45
@@
Note: We show the performance of PaECTER model in terms of Rank First Relevant (RFR), Mean Average Precision (MAP) and Mean Reciprocal Rank (MRR) at k=10, and compare the same with other models. MAP metric uses 5 positive and 25 easy negative patents. RFR and MRR@10 use just the first positive. SPECTER and SPECTER 2.0 are pre-trained models for scientific texts. PAT SPECTER is a SPECTER 2.0 model fine-tuned on our training data. PaECTER is BERT for Patents model fine-tuned on our training data. Scores are rounded off to two decimal places. Bold values represent the optimum in each column.

### v2

@@ Abstract
PaECTER is an open-source document-level encoder specific for patents. We fine-tune BERT for Patents with examiner-added citation information to generate numerical representations for patent documents. PaECTER performs better in similarity tasks than current state-of-the-art models used in the patent domain. More specifically, our model outperforms the patent specific pre-trained language model (BERT for Patents) and general-purpose text embedding models (e.g., E5, GTE, and BGE) on our patent citation prediction test dataset on different rank evaluation metrics. PaECTER predicts at least one most similar patent at a rank of 1.32 on average when compared against 25 irrelevant patents. Numerical representations generated by PaECTER from patent text can be used for downstream tasks such as classification, tracing knowledge flows, or semantic similarity search. Semantic similarity search is especially relevant in the context of prior art search for both inventors and patent examiners.
PaECTER: Patent-level Representation Learning using Citation-informed Transformers
Mainak Ghosh   Michael E. Rose   Sebastian Erhardt   Erik Buunk   Dietmar Harhoff Max Planck Institute for Innovation and Competition {mainak.ghosh, michael.rose, sebastian.erhardt, erik.buunk, dietmar.harhoff}@ip.mpg.de
@@ 1 Introduction
Domain-specific language models have become increasingly critical in the social sciences and technological domains, enabling the transformation of text into valuable data and offering scalable ways to learn semantic similarity across large document collections. In the context of patents, social scientists have shown particular interest in measuring patent similarity to analyze knowledge flows, estimate patent value, and study technological frontiers. For patent examiners, the ability to retrieve relevant prior art precisely and efficiently is essential for timely assessments of the patentability of claimed inventions under patent law and ensuring the quality of patent examinations. Missing any relevant prior patents during prior art searches would risk granting suboptimal patents, which can result in opposition or litigation, increase the burden of rework for the patent office, and undermine the objectives of the patent system.
However, existing state-of-the-art tools for document similarity learning are often specifically tailored to scientific English, which structurally and lexically differs from patent language. The PaECTER model (Patent Embeddings using Citation-informed TransformERs) addresses this gap by encoding patents into numerical representations to enable the identification of semantically similar patents.
@@
The introduction of Bi-Directional Encoder Representations from Transformers (BERT) by Devlin et al. (2019) marked a significant advancement for researchers working with text. These models, trained on large textual corpora, produce fixed-size dense vector representations (embeddings) of input text. BERT and its derivatives rely on a predefined vocabulary of tokens, which are either words or sub-words. Out-of-vocabulary tokens are either discarded or broken into smaller, recognizable subwords. This approach can result in incomplete linguistic coverage, particularly in domains with specialized language, motivating the development of language-specific BERT derivatives.
@@
Researchers have responded by developing BERT models tailored to specific domains and corpora. Among them, BERT for Patents Srebrovic and Yonamine (2020) and SciBERT Beltagy et al. (2019) are pre-trained language models optimized for patent texts and scientific language, respectively. Vocabulary selection plays a crucial role in determining the effectiveness of these models for domain-specific tasks. Our analysis highlights this by showing that only approximately 50% of the vocabulary used in SciBERT overlaps with that of BERT for Patents — underscoring the need for specialized models in patent-related research.
However, these models face a limitation when it comes to measuring text similarity. Their training objectives — predicting masked tokens and the next sentence — are not inherently designed for identifying similar documents.
To address this limitation, it is necessary to define an appropriate similarity signal and adopt a similarity-focused training objective. Contrastive learning is a machine learning approach well-suited for this purpose. It trains models to identify patterns by emphasizing the recognition of similarities and differences between data points. A key component of this method is the triplet loss function, which relies on specially crafted training data containing explicit signals for similarity and dissimilarity.
@@
SPECTER (Cohan et al., 2020) and SPECTER2 (Singh et al., 2023) exemplify the application of this approach. Both models build on SciBERT’s ability to represent scientific language and utilize training datasets composed of triplets. Each triplet includes a query, a related document (e.g., a citation), and an unrelated one (e.g., a non-citation). The original SPECTER model was trained on 684,000 triplets derived from 146,000 query documents, while the more advanced SPECTER2 was trained on over 6 million triplets. By encoding the titles and abstracts of scientific publications, both models have demonstrated superior performance on various downstream tasks in the scientific domain, including topic classification, citation prediction, and recommendation.
@@
SEARCHFORMER Vowinckel and Hähnke (2023) applies a similar approach in the patent domain. However, it is designed specifically for comparing patent claims, making it unsuitable for tasks focused on patent titles and abstracts. Additionally, the model is not publicly available.
We strive to fill a gap in the literature by providing a model that combines domain-specific vocabulary, citation-induced similarity training, and public availability. To this end, we adopt the contrastive learning technique akin to the SPECTER/SPECTER2 models. Our model is built on “BERT for Patents”, with a training objective centered on training the model to distinguish cited from non-cited patents based on semantic similarity to a focal patent. Figure 1 shows how the different models relate to each other.
Figure 1: Relationship of PaECTER model to existing models
In total, we use 300,000 patent families with a family member in the EPO, and generate five triplets for each.
In comparisons with state-of-the-art models, PaECTER shows superior capabilities in predicting semantic similarity. PaECTER supports downstream tasks in the patent domain, such as topic classification, citation prediction, and recommendation. As such, PaECTER is particularly helpful in the time-consuming patent examination task.
@@ 2 Training Data
Our training data involves 300,000 English-language patent families, including applications filed with the European Patent Office (EPO) from 1985 to 2022. A patent family refers to those patent applications and grants that originate from the same set of priority filings but seek protection in different jurisdictions.
We specifically select EPO patents from these families for two critical reasons. First, references cited in EPO search reports are under the full control of examiners, mitigating any strategic bias typically associated with citations made by inventors or patent attorneys. Second, these references are classified with a unique emphasis on their contribution to the claimed invention Webb et al. (2005), a feature not present in the United States Patent and Trademark Office (USPTO).
@@ 2.1 Objective
The primary goal of our model is to learn textual similarity and dissimilarity. By feeding the model examples of similar and dissimilar documents in relation to a focal document, the model autonomously learns optimal weights that minimize the loss function. Triplet margin loss is an effective approach in this context, where the optimizations of the weights are guided by maximizing the differences between similar and dissimilar examples.
Our training dataset comprises triplets of patents with a focal patent, a similar patent (a positive example), and a dissimilar patent (a negative example). Similar patents are those that the focal patent cites directly with a specific citation category. Dissimilar patents are those it does not cite, subject to further refinements.
The triplet margin loss function is defined as:
@@
where $\lVert.\rVert_{2}$ represents the $L_{2}$ norm distance, and $V_{F},V_{P},V_{N}$ are numerical representations for the focal patent ($P_{F}$), positive patent ($P_{P}$), and negative patent ($P_{N}$), respectively. The margin variable ($m$) ensures that the positive patent is closer to the focal patent than the negative one by at least a specified margin.
@@ 2.2 Focal Patents
We selected 300,000 patent applications as focal patents from the PATSTAT 2023 Spring version, [^1] PATSTAT is a proprietary database administered by the European Patent Office (EPO), containing bibliographical and legal event data for patents worldwide. See https://www.epo.org/en/searching-for-patents/business/patstat following specific criteria:
@@ 2.3 Positive Citations
Positive citations are sampled with replacement from the eligible pool for each focal patent, based on specific citation categories (Table 1). Of these categories, X, Y, I, and A are particularly relevant, as they indicate a strong similarity or relevance of the corresponding cited patent to the focal patent.
@@
X Particularly relevant if taken alone - prejudicing novelty.
I Particularly relevant if taken alone - prejudicing inventive step.
Y Particularly relevant if combined with another document of the same category.
A Documents defining the state of the art and not prejudicing novelty or inventive step.
O Non-written disclosure.
P Intermediate document.
T Theory or principle underlying the invention.
E Earlier patent application, but published after the filing date of the application searched (potentially conflicting patent documents).
D Document cited in the application.
L Document cited for other reasons.
& Member of the same patent family, corresponding document
Source From Webb et al. (2005). Categories in bold font are the ones we are using. Category “I” has been introduced since April 2011 (see https://register.epo.org/help?topic=citations&lng=en).
Categories X, Y, and I reflect citations where the cited patent is considered highly similar and potentially prejudicial to the novelty or inventive step of the focal patent. Although not novelty-destroying, citations of kind A reflect the state of the art for the focal patents. Accordingly, we treat cited patents of kinds X, Y, I, and A as similar to the focal patents.
@@ 2.4 Negative Citations
Negative patents, representing dissimilarity with the focal patents, are selected with a nuanced approach. First, they must come from the universe of all patents ever cited by any EPO patent, which was filed during the 1985–2022 period. Second, at least the abstract is available in English through the cited patent itself or its best substitute. Third, a CPC category must exist.
@@
  • Easy Negatives: Non-cited patents in the same CPC class and published within five years prior to the focal patent, and not indirectly cited.
@@
Figure 2 illustrates the selection process. The number of all potential easy negatives is 6,484,216; there was no focal patent that did not meet the requirements of at least three easy negatives.
@@ 2.5 Test Dataset
For the test dataset, we construct 1000 samples consisting of 1 focal patent with 5 positive citations and 25 negative citations. The negative citations are a random sample of 10 hard negatives and 15 easy negatives of a patent. Since the dataset is smaller, only focal patents were selected where there were enough positive and negative examples. We do not sample with replacement.
@@
Patent documents frequently belong to patent families, which is also the case for our test sample. 976 of our test documents belong to a multi-authority patent family. 921 of these include a US sibling. Thus our sample is not only representative for EPO patents, but also for USPTO patents.
@@ 3 Training
We build our model utilizing the sentence transformers architecture Reimers and Gurevych (2019) and Hugging Face framework Wolf et al. (2020). At its core, the model uses BERT for Patents Srebrovic and Yonamine (2020), which is connected to an input module and a subsequent document-level embedding module.
The input module concatenates a patent’s title and abstract and passes the concatenated text to the BERT for Patents model. Because of the architecture of BERT models, only the first 512 tokens (about 393 words) are used. All the 1,024-dimensional tokens of the output layer of the BERT for Patents model subsequently go to the document-level embedding module, which produces the mean embedding of all these tokens as the final numerical representation for the given input patent text. [^3] We experimented with the mean embedding of the output tokens from the BERT for Patents model; it outperformed the embedding of the [CLS] token in the similarity learning task. Hence, we do not use the [CLS] token as the final representation of any given text. During training, we feed each triplet to the model. It then optimizes the triplet loss function (Eq. (1)) by comparing the numerical representations of all three documents in that triplet.
We do minimal hyperparameter-tuning for training, following the leads of SPECTER and SEARCHFORMER. Post-tuning, we use a learning rate of 1e1e-55, 10% of the training data for warmup steps, and a triplet loss margin ($m$) of 1 for training, in addition to the Decoupled Weight Decay Regularization optimizer Loshchilov and Hutter (2019).
We train the model for four epochs, though performance plateaus after the second. Each training epoch spans approximately 20 hours, with validation occurring every 4000 steps.
We use a batch size of 4 (the maximum supported by the GPU memory) and gradient accumulation of 8 on four NVIDIA A100-SXM4-40GB GPUs for faster training and achieving an effective batch size of 128. To train the model with multiple GPUs, we have adapted the code base of the sentence transformers package slightly, and also utilized Hugging Face’s Accelerate library.
@@ 4.1 Performance Evaluation
We conduct a rank-aware evaluation of our trained model, PaECTER, and compare it against several other models and algorithms on our test dataset. The baselines of our evaluation include the BM25 algorithm, BERT, SciBERT, SPECTER, SPECTER version 2 (BASE), BERT for Patents, PatentSBERTa (Bekamiri et al., 2024), and more recent state-of-the-art, general-purpose text embedding models such as GTE (Li et al., 2023), BGE Xiao et al. (2023), and E5 (Wang et al., 2024).
We use three evaluation metrics. The first, Rank First Relevant (RFR), is the position of the first actual reference patent within the ranked list of related patents for each sample. Secondly, we use Mean Average Precision (MAP), which measures the average precision of identifying the five actual reference patents within the ranked list of 30 candidate patents for each sample, and then averages these precision scores across all samples. Finally, we use Mean Reciprocal Rank @ 10 (MRR@10), which averages the reciprocal ranks of the first correctly predicted patent among the top 10 ranked patents, averaged over all test samples.
Moreover, since pooling strategies can potentially affect model performance, we consider two widely used pooling strategies during evaluation, namely CLS and mean. Table 2 presents the performances of all models under consideration for both strategies.
In Table 2, all three metrics point to the superiority of the PaECTER model over all other models, regardless of the pooling strategy used. In particular, our trained model with mean pooling achieves higher accuracy over BERT for Patents by $7.85$ points (significant, $p<.001$) in MAP and $7.76$ points (significant, $p<.001$) in MRR@10. Furthermore, Figure 3 illustrates that, in terms of the RFR metric, our model consistently ranks the first relevant patent among the top positions compared with other models.
@@
BERT-large-uncased 2.92 1.98 38.17 51.69 58.95 71.06
SciBERT-scivocab-uncased 3.07 1.91 37.76 51.31 58.80 73.45
@@
PatentSBERTa 1.68 1.71 56.97 57.56 77.89 77.26
@@
Note: We show the performance of the PaECTER model in terms of Rank First Relevant (RFR), Mean Average Precision (MAP), and Mean Reciprocal Rank (MRR) at $k=10$, and compare it with other models. The MAP metric uses 5 positive and 25 negative patents (10 hard negatives and 15 easy negatives). RFR and MRR@10 use only the first positive. Scores are rounded to two decimal places. Bold values represent the optimum in each column.
@@
4.2 Comparison with SEARCHFORMER
SEARCHFORMER is another state-of-the-art model specifically designed for patent examination by Vowinckel and Hähnke (2023). It was trained specifically to assist EPO patent examiners in prior art search. Unlike PaECTER, it was trained on claims texts rather than titles and abstracts. SEARCHFORMER is a closed-source language model and the training dataset is confidential.
To assess the relative performance of PaECTER, we compare its rank predictions with those of SEARCHFORMER using SEARCHFORMER’s test dataset. The goal of this evaluation is to suggest relevant prior art given the first claim of a patent.
From the authors, we obtained about 90% of the original SEARCHFORMER test dataset, excluding pending applications (which are confidential). This subset includes 1,795 focal patents and a total of 1,154,557 pairs. Each focal patent is associated with 1 to 14 positive citations and between 197 and 955 negative citations.
Table 3: Rank-aware evaluation of PaECTER against SEARCHFORMER using the SEARCHFORMER dataset
Model Avg. RFR MAP MRR@10
SEARCH FORMER 58.73 9.38 15.33
PaECTER 49.66 11.13 17.93
Note: We show the performance of the SEARCHFORMER and PaECTER models in terms of Rank First Relevant (RFR), Mean Average Precision (MAP), and Mean Reciprocal Rank (MRR) at $k=10$. The MAP metric uses 5 positive and 25 negative patents (10 hard negatives and 15 easy negatives). RFR and MRR@10 use only the first positive. Scores are rounded to two decimal places. Bold values represent the optimum in each column.
Table 3 portrays the results of this exercise. It shows that PaECTER outperforms SEARCHFORMER in all metrics. Based on the RFR metric, we see that on average, an examiner would have to inspect nearly 50 documents suggested by PaECTER to find relevant prior art, which is 9 documents fewer than with SEARCHFORMER.
4.3 External Evaluation
Ganguli et al. (2024) have tested the PaECTER, among other models, on several tasks involving independent claims of US patents published between 1850 and 2023. Other models include S-BERT and GTE, next to a TF-IDF representation as a baseline model. Ganguli et al. (2024) compare these models in three tasks: (i) an interference task, (ii) a human annotation task, and (iii) a technology classification task.
The interference task was to connect two (pending) patents on the same invention by two independent inventors. The task used decisions were made by patent examiners between 1998 and 2014. Results show that PaECTER outperforms all other models, though GTE follows second by a small margin.
In the human annotation task, the models had to replicate non-expert human judgment on which patent was more similar in a pair-wise comparison. The task included patents published during the 1880–1920 period. In this task, PaECTER comes third with GTE being best. As the authors note, this is likely due to the historical data involved in the test.
Finally, the technology classification task involved predicting whether a pair of patents belongs to the same CPC class (9 in total) or CPC section (123 in total). The test included patents published from the 1850–2023 period. Depending on the metric (ROC AUC or PR AUC), PaECTER ranks first or second (behind S-BERT) in predicting whether two patents belong to the same CPC class. It ranks second or third in predicting shared CPC section membership, though all models perform similarly on this task.
4.4 Ablation method
We perform an ablation study whereby we include only title and CPC classes, but not abstracts, in the training of the model. The Cooperative Patent Classification (CPC) is the most used technology classification system and available for almost every patent (Lobo and Strumsky, 2019). It consists of sections, classes, subclasses, main groups and subgroups, thereby providing a hierarchical or nested classification. At the lowest granularity, there are well more than 250k CPC subgroups. Most patents belong to several CPC subgroups (Lobo and Strumsky (2019) found an average of 5 for patents granted by the USPTO between 1975 and 2018).
Since not every patent has CPC codes, the training, validation and test sets decrease somewhat in size as compared to the original application. For example, about 103k patents from the original training set lack CPC codes.
Table 4: Ablation method with CPC codes instead of abstracts
Model/Test Avg. RFR MAP MRR@10
CLS Mean CLS Mean CLS Mean
Original Model / Original Testset 1.32 1.31 66.91 68.17 88.03 88.25
Original Model / Ablated Testset 2.01 1.92 56.59 57.63 71.87 72.1
Ablated Model / Original Testset 1.59 1.48 57.37 63.59 79.23 82.47
Ablated Model / Ablated Testset 1.88 1.83 59.4 60.55 73.96 75.04
Note: We show the performance of the PaECTER model and the ablated PaECTER model in terms of Rank First Relevant (RFR), Mean Average Precision (MAP), and Mean Reciprocal Rank (MRR) at $k=10$. The MAP metric uses 5 positive and 25 negative patents (10 hard negatives and 15 easy negatives). RFR and MRR@10 use only the first positive. Scores are rounded to two decimal places. Bold values represent the optimum in each column. The ablated test set has about 5k samples less than the original test set.
Table 4 presents the results. We perform the test separately for training data with CPC codes instead of abstracts, as well as for ablated test set. For all the metrics and pooling methods we find that the original model (trained on abstracts) outperforms the ablated model.